Non-exercise activity thermogenesis (NEAT) provides a partial explanation for the previous result. NEAT is energy burned by fidgeting, maintenance of posture, spontaneous muscle contraction, and other activities of daily living. In the study of the previous exercise, the 16 subjects increased their NEAT by 328 calories per day in response to their additional food intake. The standard deviation was 256. (a) Test the null that there was no increase in NEAT versus the 2 sided alternatives. Summarize the results and state your conclusion. (b) Find a 95% CI for the change in NEAT. Discuss the additional information provided by the CI that is not evident from the results of the significance test.
Answer #1

Given X = 328 . s = 256 n =16 Test hypothesis are Null hypothesis Ho: u=0 Alternate hypothesis Hau #0 Test statistic Test sta

b)

95% confidence interval is : Critical t df=n-1 =15 Level of confidence C = 0.95 a = 0.05 Critical value tc=ta/2.df = 2.131 (t